# A General Theory of Abiogenesis: Thermodynamic Selection and the Emergence of Life

### v5

@@ 9 Model Flexibility and Testing
These qualities make the framework extensible. While simple in formulation, it is meant as a conceptual scaffold capable of supporting more detailed physical and chemical modeling in future work.
Test case scenario: Entropy-Driven Self-Assembly in Amphiphilic Systems
A tractable and highly illustrative system for testing TALM involves the self-assembly behavior of amphiphilic molecules in water, where structure formation is driven primarily by the hydrophobic effect. This effect arises because water molecules form hydrogen-bonded networks. When non-polar (hydrophobic) molecules are introduced, they disrupt these networks, lowering the system’s entropy. Water responds by ordering tightly around the intruding molecules. However, when many hydrophobic molecules are present, they tend to cluster together, freeing the surrounding water to return to a disordered state, thus increasing entropy.
To explore this within the TALM framework, consider a series of aqueous systems each containing 18 mL of water (1 mole) and 0.01 mole of one of three solutes: sodium acetate, sodium octanoate, and sodium dodecanoate. These molecules differ only in the length of their hydrophobic tails. When air is gently introduced, observable differences arise. Sodium acetate, which lacks a hydrophobic tail, dissolves completely with no visible change. Sodium octanoate, with a moderate tail length, forms a transient bubble. Sodium dodecanoate, with a longer hydrophobic chain, forms a stable bubble and, under microscopic inspection, many vesicles.
These behaviors are predicted by TALM as a function of chain length and the system’s ability to manage entropy and energy flow. The longer tails drive stronger self-association, lowering diffusion ($D(x,t)$) and increasing structural resilience ($R_{n}$), while also allowing water to release more entropy as structured cages collapse. Thus, the entropy cost of forming vesicles is more than offset by entropy export to the environment:
$\Delta S_{net}=\Delta S_{env}-\Delta S_{local}>0$
leading to a reduced $\Phi(t)$ and positive $y^{\prime}(t)$.
TALM predicts that structural persistence-bubbles, micelles, and vesicles-emerges at a critical chain length $L_{c}$. Based on this system:
  • For $L<C8$, self-assembly is minimal; no structures persist.
  • At $L\approx C10$, vesicle formation becomes likely.
  • Between $C10$ and $C14$, vesicles are increasingly stable.
  • Beyond $C14$, solubility may decrease, limiting structure formation.
This example shows how TALM can be experimentally tested using simple compounds and observable phenomena. It connects entropy regulation, diffusion control, and structural stability into a single framework and reinforces TALM’s core claim: persistence arises where entropy export aligns with structural coherence under fluctuating energy.
10 Theoretical Outlook
The TALM framework provides a means of translating this conceptual model into a probabilistic structure applicable across diverse environmental conditions. The reaction viability inequality, extended via $y^{\prime}(t)$, allows for the estimation of persistence under spatial and energetic modulation. This yields not only a theoretical basis for selection-like behavior in prebiotic chemistry but also a potential comparative lens across planetary systems.
Continued refinement of this approach, through simulation of reaction networks, laboratory analogs of compartmentalized cycling systems, and the thermodynamic profiling of planetary bodies, may illuminate the constraints and pathways most conducive to the emergence of persistent reaction architectures. In this light, the TALM framework serves not only as a theoretical advance but as a scaffold for the systematic comparison of environments where the thermodynamic roots of natural selection may have taken hold.
Moreover, pairing the TALM framework with Urability, the environmental pre-screening framework (Deamer, 2021), which assesses whether a planetary setting possesses the thermodynamic and chemical conditions necessary for prebiotic complexity, provides a two-tiered approach by first evaluating urability, then applying TALM. This provides a robust, scalable method for identifying environments most conducive to abiogenesis.
Universal Scope and Molecular Agnosticism
@@ 10 Theoretical Outlook
By grounding the emergence of selection-like behavior in thermodynamic alignment rather than chemical specificity, TALM invites the consideration of exotic life forms, composed of silicon, metal-surface catalysis, quantum structures, or unknown substrates, as valid targets for origin-of-life investigation. The only constraint imposed is the existence of fluctuating energy, compartmentalizing structure, and the potential for local entropy reduction consistent with overall entropy production. This approach is consistent with the theoretical perspectives of Davies and Cleland, who argue that biology as currently known may only reflect a narrow subset of possible life-supporting structures (Cleland and Chyba, 2002; Davies, 2009).
@@
allows for extrapolation from localized chemical systems ($N$ as individual reaction permutations) to whole-planetary or even galactic estimations ($N$ as the total number of habitable micro-environments or reaction domains). This scaling supports the view that abiogenesis may be not merely possible but statistically probable across large enough thermodynamic landscapes. Similar views have been proposed in the context of cosmological selection (Lineweaver et al., 2002; Walker et al., 2016), where life is treated as a common feature of entropy-utilizing complexity.
In this light, TALM does not model the origin of “life” in a strict biological sense, but rather the origin of persistent structure, thermodynamic systems that exhibit selective retention within entropic environments. It therefore provides a potential bridge between prebiotic chemistry, astrobiology, and universal principles of organized complexity.
Implications for Abiogenesis
@@
  • Reaction networks that dissipate energy effectively (England, 2013)
  • Autocatalytic or structurally recursive sets (Kauffman, 1993)
  • Systems capable of storing energy or buffering against loss (Pross, 2012)
  • Structures arising from entropy-compatible self-assembly (Prigogine and Nicolis, 1977; Schneider & Sagan, 2005)
@@
Consequently, the TALM framework, by treating life emergence as a thermodynamic selection process, leads to several key predictive implications:
  • Life-like systems are statistically inevitable in diverse environments where fluctuating energy and chemical complexity coexist (England, 2013).
  • Non-replicative prebiotic systems may be far more common than those capable of heredity (Walker and Davies, 2013).
  • Gas giants, icy moons, and subsurface oceans could all plausibly host persistence-based chemical systems, even in the absence of traditional biosignatures (Seager, 2013).
  • Life emergence may occur in multiple distinct epochs on a single body, driven by cyclical thermodynamic phases rather than one-time events (Bains, 2004).
  • Exoplanetary systems will exhibit structure favoring thermodynamically stable bodies, potentially leading to testable correlations between exoplanet class and biopotential (Seager, 2013).
@@ 11 Conclusion
The predictive application of TALM to an aqueous amphiphilic system illustrates the theory’s utility beyond abstract thermodynamic reasoning. By focusing on a simple comparative setup-acetate, octanoate, and dodecanoate under identical water and energy conditions, we show that TALM can forecast the onset of vesicle formation as a function of chain length. This experiment highlights a specific thermodynamic threshold, shaped by entropy export and diffusion constraints, where spontaneous structural persistence begins to emerge. [^1] See Appendix C for additional proposed experimental systems designed to explore structure-environment compatibility under diverse thermodynamic constraints. These findings align with classical hydrophobic theory while extending it through a persistence-focused framework.
We have presented a thermodynamic model that reframes the emergence of natural selection as a consequence of energy alignment in prebiotic reaction networks. Grounded in a reaction viability inequality, the framework captures how fluctuating environmental energy can drive the persistence, or collapse, of molecular systems, independent of heredity or replication.
By formalizing this principle with defined parameters for energy requirements ($x_{i}$), releases ($r_{i}$), storage ($s_{i}$), and network resilience ($R_{n}$), we demonstrate that differential persistence can occur even in non-replicative systems. Such a filter offers a pre-Darwinian mechanism of selection, one that operates purely through energy alignment, without heredity or reproduction.
The model is contextualized within foundational work by Kauffman, Prigogine, Deavey, England, and Pross on dissipative structures, organizational closure, and self-organization and then extended to include entropy production, diffusion constraints, and spatial modulation, enriching the model’s physical plausibility and bridging it to realistic prebiotic environments.
The Thermodynamic Abiogenesis Likelihood Model (TALM) expands this core framework to a planetary scale, allowing us to estimate the probability of persistent chemical systems arising under specific energy cycling and spatial conditions. By defining parameter ranges and comparative planetary contexts, TALM opens new avenues for assessing the likelihood of abiogenesis across a spectrum of celestial bodies.
This expansion enriches the model without compromising its principle elegance, and has a workable scope in its usage in future developments through simulation, environmental specificity, and probabilistic modeling across interstellar system settings. The incorporation of entropy and compartmentalization strengthens the physical plausibility of early self-organizing systems and when combined with environmental filters such as urability, positions the framework as a modular and physically grounded strategy for assessing the likelihood of life’s emergence across diverse planetary scenarios.
@@
Crucially, TALM is both scalable and agnostic to molecular identity. It applies not only to terrestrial carbon-based chemistry but to any structured environment where energy gradients and entropy dynamics permit the formation of persistent configurations. Its scaling from local to galactic domains suggests that the emergence of life-like systems may be a statistical outcome of thermodynamic diversity across the universe.
This model invites a reconsideration of abiogenesis not as a rare biochemical accident, but as a recurring feature of structured, energy-driven systems. Selection may begin not with genes, but with thermodynamic compatibility, an emergent signature of matter shaped by energy and constrained by entropy.
In conclusion, we propose that selection begins not with genes, but with energy. This framework offers what may be the most physically grounded resolution yet to one of science’s most enduring mysteries: how life emerges from non-life. By identifying thermodynamic compatibility, not molecular instruction, as the initial driver of persistence, this model reframes abiogenesis as a natural consequence of energy flow, entropy modulation, and structured environments.
Appendix A: Model Parameter Ranges
Parameter Description Typical Range
$x_{i}$ Energy required per reaction $10^{-21}$ to $10^{-18}$ J
$r_{i}$ Energy released per reaction $10^{-21}$ to $10^{-18}$ J
$s_{i}$ Stored energy (buffer) Up to $r_{i}$
$R_{n}$ Network resilience 0–1 (normalized)
$\Delta S_{net}$ Net entropy balance $10^{-22}$ to $10^{-20}$ J/K
$D(x,t)$ Diffusion coefficient $10^{-14}$ to $10^{-9}$ m2/s
$z(t)$ Environmental input Sinusoidal, stochastic, or empirical
Table 3: Proposed parameter ranges for TALM model components.
Appendix B: Alternate Persistence Regimes - Experimental Proposals
TALM predicts that viable, persistent structures should emerge in systems where entropy export, energy input, and diffusion constraints are properly balanced. Below are several proposed experimental regimes to test this hypothesis beyond the aqueous amphiphile scenario:
B1. Thermal Cycling in Viscous Organic Media
  • Hypothesis: Viscosity and solvent polarity affect molecular organization and persistence.
  • Environment: Ethylene glycol or glycerol with amphiphilic or prebiotic molecules, cycled between 20–80C.
  • Prediction: Micelle-like or polymerizing structures may appear and persist through multiple cycles.
B2. Non-Aqueous Solvent Systems
  • Hypothesis: Alternative solvents may support reverse micelles or other non-lipid-based order.
  • Environment: Formamide, ethanol, or water-formamide mixtures with organic acids.
  • Prediction: Stable clusters or scaffolds may form that persist despite solvent polarity changes.
B3. Wet-Dry Cycling with Polymer Precursors
  • Hypothesis: Dehydration increases entropy flux and favors polymer formation.
  • Environment: Thin film surface hydration-dehydration using amino acids or hydroxy acids.
  • Prediction: Polymer chains persist across cycles and exhibit increased resistance to hydrolysis.
B4. Freeze-Thaw Cycles in Brine Ice Matrices
  • Hypothesis: Icy matrices may stabilize fragile structures by limiting diffusion and allowing energy buildup.
  • Environment: Slow freeze-thaw cycles in saline water or formamide solutions.
  • Prediction: Weakly bound polymers or colloids may persist and restructure after thawing.
Appendix C: Test Case Scenarios for TALM Validation
Table 4: Candidate experimental scenarios to test TALM predictions in prebiotic-like systems. Each setup is designed to probe specific parameters in the Reaction Viability Inequality.
Scenario ID Description Testable Variable(s) Prediction Based on TALM Suggested Methodology
C1 Vesicle persistence under thermocycling (e.g., sodium acetate vs. dodecanoate) Chain length, entropy gradient ($\Delta S$), hydration cycle rate Longer chains with moderate amphipathic properties will persist more stably in fluctuating entropy environments Thermocycling wet-dry system with microscopy or turbidity assays
C2 Stability of co-polymers with differing energy costs per bond Bond energy ($\Delta E$), cycle time ($\tau$) Systems with intermediate energy bonds optimized for available $\Delta E/\tau$ will show maximum persistence Synthetic polymerization assays with calorimetric energy profiling
C3 Competitive selection of reaction networks in microfluidic gradient landscapes Flow-driven $\Delta S$, competing chemical pathways More entropy-efficient or energy-compatible reactions will dominate spatially Microfluidic droplet networks with selective tracing (e.g., fluorescence)
C4 Entropy-driven assembly of amphiphilic structures Lipid concentration, gradient steepness System will preferentially select configurations with lowest dissipation per persistence unit Vesicle/micelle quantification under controlled heating gradients
C5 Chain length vs. structural stability in varying environmental entropy rates Chain length ($n$), external $\Delta S$ rate There exists an optimal range where persistence probability peaks (as per $Z>0$) Incremental chain length synthesis with stability and lifespan tracking
C6 Reversible vs. irreversible reaction dynamics in cycling environments Reaction reversibility, $\tau$, entropy change Irreversible reactions with moderate $\tau$ and favorable $\Delta E$ will outcompete reversible ones in persistence likelihood Isothermal titration calorimetry + cyclic environment simulation
Acknowledgements
@@
The author acknowledges the use of generative AI tools during the drafting and editing stages of this manuscript, particularly for assistance with structuring, clarity, and technical expression. All scientific reasoning, theoretical development, and interpretation remain the responsibility of the author.

### v7

@@ 8.6 Scientific risk
The test case scenarios proposed here, show how the model can be experimentally tested using simple compounds and observable phenomena. They connect entropy regulation, diffusion control, and structural stability into a single framework and reinforces the core claim: persistence arises where entropy export aligns with structural coherence under fluctuating energy.
8.7 Universal Scope and Molecular Agnosticism
@@
By grounding the emergence of selection-like behavior in thermodynamic alignment rather than chemical specificity, TALM invites the consideration of exotic life forms, composed of silicon, metal-surface catalysis, quantum structures, or unknown substrates, as valid targets for origin-of-life investigation. The only constraint imposed is the existence of fluctuating energy, compartmentalizing structure, and the potential for local entropy reduction consistent with overall entropy production. This approach is consistent with the theoretical perspectives of Davies and Cleland, who argue that biology as currently known may only reflect a narrow subset of possible life-supporting structures [31, 32].
@@
allows for extrapolation from localized chemical systems ($N$ as individual reaction permutations) to whole-planetary or even galactic estimations ($N$ as the total number of habitable micro-environments or reaction domains). This scaling supports the view that abiogenesis may be not merely possible but statistically probable across large enough thermodynamic landscapes. Similar views have been proposed in the context of cosmological selection [33, 34], where life is treated as a common feature of entropy-utilizing complexity.
In this light, the Thermodynamic Abiogenesis Likelihood Model does not model the origin of “life” in a strict biological sense, but rather the origin of persistent structures where thermodynamic systems exhibit selective retention within entropic environments. It therefore provides a potential bridge between prebiotic chemistry, astrobiology, and universal principles of organized complexity.
8.8 Implications for Abiogenesis
@@
  • Reaction networks that dissipate energy effectively [2]
  • Autocatalytic or structurally recursive sets [1]
  • Systems capable of storing energy or buffering against loss [3]
  • Structures arising from entropy-compatible self-assembly [12, 11]
@@
Consequently, the Thermodynamic Abiogenesis Likelihood Model, by treating life emergence as a thermodynamic selection process, leads to several key predictive implications:
  • Life-like systems are statistically inevitable in diverse environments where fluctuating energy and chemical complexity coexist [2].
  • Non-replicative prebiotic systems may be far more common than those capable of heredity [35].
  • Gas giants, icy moons, and subsurface oceans could all plausibly host persistence-based chemical systems, even in the absence of traditional biosignatures [36].
  • Life emergence may occur in multiple distinct epochs on a single body, driven by cyclical thermodynamic phases rather than one-time events [37].
  • Exoplanetary systems will exhibit structure favoring thermodynamically stable bodies, potentially leading to testable correlations between exoplanet class and biopotential [36].
@@
The theoretical outlook presented here positions the Thermodynamic Abiogenesis Likelihood Model as both a bridge and a filter between conceptual thermodynamics and experimental reality. By defining persistence in measurable energetic terms, the model converts the question of life’s emergence from one of abstract plausibility into one of empirical exposure. Each proposed test therefore represents a deliberate opportunity for the framework to fail - a necessary criterion for genuine scientific progress. If the predicted signatures of persistence are not observed, the theory must be revised or rejected; if they are, then the principle that thermodynamic alignment precedes biological heredity gains tangible support. In this sense, tests are specifically designed to satisfy the criterion that a sound theoretical model must not only describe what might occur but also specify the precise conditions under which it would not.
@@ 9 Conclusion
The predictive application of the persistence inequality to aqueous amphiphilic systems demonstrates the model’s utility beyond abstract thermodynamic reasoning. By examining acetate, octanoate, and dodecanoate under identical conditions, we identify a chain-length threshold at which vesicle formation becomes thermodynamically favored. This aligns with established hydrophobic theory while reframing it through the lens of energetic persistence. It also provides a concrete and falsifiable prediction: when energy intervals and chain length surpass defined thresholds, spontaneous structural organization should appear.
Grounded in a reaction viability inequality, the framework explains how fluctuating environmental energy can drive the persistence or collapse of non-replicative molecular systems. By formalizing energy requirements ($x_{i}$), releases ($r_{i}$), storage ($s_{i}$), and resilience ($R_{n}$), it shows how differential persistence can arise in the absence of heredity or replication, offering a genuinely prebiotic mechanism of selection. In this view, selection emerges as a thermodynamic property of open systems rather than a biological innovation.
Contextualized within the foundational work of Kauffman, Prigogine, Deamer, England, and Pross on dissipative structures and self-organization, the model integrates entropy production, diffusion constraints, and spatial modulation into a physically coherent description of early prebiotic environments. The Thermodynamic Abiogenesis Likelihood Model (TALM) extends this principle to planetary scales, enabling probabilistic assessments of persistence under specific energy-cycling and spatial conditions and providing a structured way to compare the abiogenic potential of different worlds.
@@
Crucially, TALM is scalable and agnostic to molecular identity. Its criteria apply not only to terrestrial carbon chemistry but to any structured environment where energy gradients and entropy dynamics permit the formation of persistent configurations. This broadens the plausible scope of life’s emergence across diverse planetary scenarios and challenges the assumption that abiogenesis requires rare or narrowly defined conditions.
If the predictions outlined here are verified, the origin of life is no longer an improbable chemical accident but a measurable physical outcome of energy/entropy alignment. The proposed experimental scenarios translate the abstract viability condition into operational tests, linking the persistence inequality to known systems such as micelles, folding peptides, and surface-assisted chemistry. Combined with the Thermodynamic Abiogenesis Likelihood Framework, they establish a practical, physically grounded approach for identifying prebiotically favored systems - structures selected not for biological function, but for energetic persistence.
[…]
*Supplementary information
[…]
*Acknowledgments
@@
10 Declarations
  • Funding: Not applicable.
  • Conflict of interest: The author declares none.
  • Ethics approval and consent to participate: Not applicable.
  • Consent for publication: Not applicable.
  • Data availability: All data are included within the article and its supplementary information.
  • Materials availability: Not applicable.
  • Code availability: Not applicable.
  • Author contribution: T.M.Prosser is the sole author of this work.